2(x+3)=4
3(x+2)=3
4)x+3)=4
12=3(2-y)
4(x+3)=8
3(x+5)=12
4(2x+8)=88
3(5a-3)=6
-3(x-5)=-21
-3(4-y)=-27

Answer by queenofit(120)

This is simple distrubuting. YOu must multiply the number before the ( ) to each number in the ( ). then solve for the variable. I will not do all your home work , but I will show some of your problems as examples. Then you should be able to to the work... Sound fair? you need to practice.
First, always remember to watch your signs when you multiply or divide, ok.
lets do
2(x+3)=4 first multiply the 2*x and 2*3
2x+6=4 then solve for x
2x= -6+4 move variables and numbers to different sides
2x=-2 divide the number in front of variable into both sides
x=-1 watch your signs
next one lets do
-3(x-5)=-21
-3x + 15 = -21
-3x= -15-21
-3x=-36
x=12